OpenStudy (darkbluechocobo):

Assume the following angle is in standard position. Find a positive and negative angle whose absolute value is less than 360° and is coterminal with the given angle: -485°

OpenStudy (darkbluechocobo):

@amistre64

OpenStudy (darkbluechocobo):

would it look like this -485-360=-845 -485-720=-1205